OpenStudy (anonymous):

what is 2(x-5)+3(5-x)=-2?

OpenStudy (oaktree):

2x-10+15-3x=-2, so -x+5=-2, so x=-3.

OpenStudy (oaktree):

I mean x=7.

OpenStudy (anonymous):

2(x-5)+3(5-x)=-2 2x - 10 + 15 - 3x = -2 -x + 5 = -2 -x = -7 x = 7

OpenStudy (oaktree):

Oops!

OpenStudy (anonymous):

thank you hairtub

OpenStudy (anonymous):

No problem :)